Based on a recent study, roughly 80% of college students in the U.S. own a smartphone. Is the proportion of smartphone owners lower at this university?

Proportion of students who have a smartphone 78% (243 students) students with no smartphone 21% (67 students)

  1. State your hypotheses in symbolic form and in words.
  2. use a normal model to estimate the P-value probability. Verify that normality conditions are met.
  3. conduct the hypothesis test.
  4. Give your P-value and interpret its meaning as a conditional probability.
  5. State a conclusion that answers the research question. Use a significance level of 5%.
Answer #1

Here we have a P20080, 5= 243+67=310, ²20.78 *20.05 Here np.= 0, so x310= 298 hp = 0. 20 x 300 - 6 2 57 ор с 5 aud 79, s“. hеNow pualne at a = -0.88 pualues PLZ>|-0.881) 1- P125 0.88) 31-0.8106 -0.1894 since pualue 7000s, we failed to redect to concl